# SentenceKV: Efficient LLM Inference via Sentence-Level Semantic KV Caching

### v1

@@ Abstract
Large language models face significant computational and memory challenges when processing long contexts. During inference, efficient management of the key-value (KV) cache, which stores intermediate activations for autoregressive generation, is critical to reducing memory overhead and improving computational efficiency. Traditional token-level efficient KV caching methods overlook semantic information, treating tokens independently without considering their semantic relationships. Meanwhile, existing semantic-preserving KV cache management approaches often suffer from substantial memory usage and high time-to-first-token. To address these limitations, we propose SentenceKV, a novel sentence-level semantic KV caching approach designed to enhance inference efficiency while preserving semantic coherence. During prefilling, SentenceKV groups tokens based on sentence-level semantic similarity, compressing sentence representations into concise semantic vectors stored directly on the GPU, while individual KV pairs are offloaded to CPU. During decoding, SentenceKV generates tokens by selectively retrieving semantically relevant sentence-level KV entries, leveraging the semantic similarity between the prefilling-stage semantic vectors and decoding-stage queries. This ensures efficient and contextually accurate predictions, minimizing the loading of redundant or irrelevant data into GPU memory and significantly reducing memory overhead while maintaining stable inference latency, even for extremely long contexts. Extensive evaluations on benchmarks including PG-19, LongBench, and Needle-In-A-Haystack demonstrate that SentenceKV significantly outperforms state-of-the-art methods in both efficiency and memory usage, without compromising model accuracy.
@@ 1 Introduction
Large language models (LLMs) have achieved remarkable success across a wide range of natural language processing tasks [19, 30]. As LLMs are increasingly applied to complex reasoning and multi-step decision-making scenarios, the demand for longer context windows has grown significantly [6]. Recent advancements have extended the context window from the early GPT models to the latest iterations, such as GPT-o1 and beyond [20, 13]. However, this growth introduces substantial challenges in key-value (KV) cache efficiency, leading to memory bottlenecks and increased inference latency [23].
@@
Due to the transformer-based architecture of LLMs [26], the memory footprint of KV cache grows linearly with the context length, directly impacting GPU usage. For example, in the case of the LLaMA-3.1-8B model [1], as shown in Appendix A.1, processing a 32k-token prompt during the decoding stage requires approximately 16 GB (using float16 precision) of GPU memory, which can be prohibitive for users without high-end hardware. Furthermore, longer prompts significantly increase computational overhead, as each additional token requires computing attention weights across the entire sequence. This quadratic attention complexity leads to substantial compute inefficiency, making efficient inference particularly challenging in resource-constrained environments.
@@
To mitigate these issues, several methods have been proposed to compress KV cache [23]. Among them, token eviction strategies based on fixed selection patterns or attention-weight-based pruning have shown promise [27, 29]. However, these approaches suffer from key limitations: they fail to adapt to the dynamic nature of token importance during decoding and often discard tokens permanently, neglecting their potential relevance in the future decoding steps. This static behavior can degrade the performance of the model, especially in tasks requiring complex reasoning and long-range dependencies [16].
@@
Another category of approaches, such as Quest [25] and ShadowKV [24], dynamically retrieve portions of KV cache from CPU-offloaded storage back to GPU memory during attention calculation. Typically, these methods segment the input into fixed-size chunks and selectively retrieve them based on query relevance at each decoding step. However, fixed-size chunking can disrupt semantic coherence by arbitrarily breaking natural semantic boundaries. Some dynamic retrieval approaches, including ClusterKV [15] and SqueezeAttention [11], mitigate this issue by clustering semantically related keys into coherent chunks. Yet, these clustering-based methods introduce significant computational overhead, resulting in increased time-to-first-token (TTFT) latency.
@@
To address these limitations, we propose SentenceKV, a novel approach that dynamically manages KV cache based on sentence-level semantic information. Our key insight is that sentences inherently encapsulate richer semantic information compared to isolated tokens, making them a more effective unit for caching and retrieval. For instance, as illustrated in Figure 1, sentences from different domains tend to form distinct semantic clusters within the representation space of LLMs, a phenomenon also observed in recent literature [5]. We further propose a novel multi-query-aware mechanism for retrieving relevant sentence-level KV cache entries during decoding.
@@
Extensive experiments on long-context benchmarks, including LongBench [2], and Needle-In-A-Haystack (NIAH) [12] demonstrate that our approach not only improves inference efficiency but also maintains or even enhances model accuracy across diverse tasks. Quantitative analysis shows that our method reduces memory consumption by 33% compared to standard full KV cache implementations when processing context lengths up to 32k tokens, offering an efficient solution for long-context processing.
@@ 2 Related Work
StreamingLLM [27] enables infinite-context handling via attention sinks and a moving window, allowing models to process extremely long sequences without growing memory footprints. SnapKV [14] preserves essential KV positions based on per-head attention patterns with an observation window. FastGen [9] progressively drops tokens with low future impact, enhancing efficiency with minimal quality loss. PyramidInfer [28] and PyramidKV [4] propose a layer-wise pyramidal KV cache retention pattern, preserving more tokens in upper layers of the model to maximize memory savings without sacrificing context.
@@
Dynamic access mechanisms adaptively manage KV retention based on task-specific requirements. H2O [29] dynamically retains high-impact tokens based on accumulated attention scores. TOVA [21] prunes tokens based on real-time query relevance, but risks removing context needed for future steps. The quest [25] partitions KV cache into pageable memory and predicts critical pages per query, in which case fragmentation issues can arise. ShadowKV [24] offloads the value cache to CPU and stores a low-rank key cache on GPU, using a KV selection strategy to reduce memory footprint and maintain high-throughput decoding. SqueezeAttention [11] optimizes cache allocation across both sequence and layer dimensions, assigning larger budgets to important layers while pruning unimportant ones. DynamicKV [31] implements task-aware adaptive retention, adjusting KV budgets per layer to maintain task-specific relevance while reducing cache size.
@@
As LLMs tackle increasingly complex tasks, optimizing KV cache at a semantic level is crucial to maintaining output coherence. ChunkKV [17] groups tokens into semantic chunks, retaining the most informative segments and discarding redundant ones to enhance long-context inference efficiency. ClusterKV [15] introduces semantic clustering, which, unlike methods that permanently evict tokens or recall them based solely on textual positions, clusters tokens semantically and recalls them at the granularity of semantic clusters. Task-KV [10] differentiates between heterogeneous and general aggregation attention heads, preserving full KV cache for critical heads while reducing it for less relevant ones. SepLLM [5] introduces a data-dependent sparse attention mechanism to mitigate the excessive attention allocated to seemingly uninformative separator tokens. By compressing segment-level information into these tokens, SepLLM reduces KV cache while preserving essential content.
While chunk-based compression better preserves semantics than token-wise eviction, it may still fragment full thought units, omitting crucial context. Prior studies show that ignoring natural boundaries, such as sentence breaks, can harm language modeling performance [7, 22]. SentenceKV addresses this by retaining entire sentences or semantically self-contained segments, ensuring holistic context preservation. By bridging local semantic retention from ChunkKV with global coherence, SentenceKV further enhances long-context inference while maintaining minimal memory overhead.
@@ 3 Motivation
Existing KV cache compression methods typically operate at the token-level, where tokens considered less important—often determined by heuristics such as attention scores—are evicted during the prefilling stage. However, recent studies suggest that the relevance of prefilled tokens is not static throughout the decoding process [25]. Evicting tokens prematurely based on their initial perceived importance can degrade inference accuracy since tokens initially deemed irrelevant might later become crucial during decoding. To address this dynamic nature, recent studies propose retrieving essential prefilled tokens dynamically based on their relevance to the current decoding query. To mitigate computational overhead associated with searching across all prefilled tokens, these methods typically segment the prefilling text into fixed-size chunks, retrieving and utilizing only the most relevant chunks during decoding.
Figure 1: Similarities between sentence embeddings from different categories in the Pile datasets [8].
However, these chunk-based strategies lack a principled justification for preserving semantic coherence, as fixed-size chunks often fragment semantic content arbitrarily. Intuitively, sentences encapsulate complete semantic units, making them natural candidates for maintaining context coherence. To empirically validate this intuition, we examined sentences from distinct categories in the Pile dataset [8] and measured their embedding similarities in LLMs. As illustrated in Figure 1, sentences from different semantic categories exhibit clearly distinguishable embedding patterns. Moreover, during the decoding phase, we observed that the model-generated query tends to focus attention primarily on semantically related sentences rather than arbitrary token segments. This observation motivates our sentence-level caching approach, which segments the prefilling text into semantically coherent sentence units. During decoding, we leverage the sentence currently being generated to dynamically identify and recall the most semantically relevant sentence-level cache entries, thereby efficiently preserving context and enhancing inference accuracy.
@@ 4.1 Prefilling Stage
We first split the input text into sentences according to punctuation, creating multiple sentence buckets, each containing a varying number of tokens. For example, a 32K token input might be split into hundreds of sentence buckets of different lengths. While token importance changes dynamically during decoding, maintaining all tokens at full precision in GPU memory is inefficient and unnecessary. To determine which tokens to retain, we apply an observation window [14] for token importance calculation as follows:
@@
For an input prompt of length $L$ (e.g., $L=32K$), we designate the last $N$ tokens (typically $N=32$) as our observation window. We perform a forward pass where each token in positions $(L-N+1)$ to $L$ calculates attention scores with respect to all preceding tokens in positions $1$ to $(L-N)$. For each token $i$ in the preceding positions, we calculate its importance score $\alpha_{i}$ by summing the attention scores it receives from all tokens in the observation window, across all the attention heads.
@@
We set a token budget $\tau$ (e.g., $\tau=1024$) that represents the maximum number of tokens we can keep in GPU memory during decoding. Since tokens that appear unimportant during prefilling may become relevant at later decoding steps, we introduce the semantic keeping factor $r$ (typically $r=2$ or $3$), which controls how many tokens are initially retained before selective pruning during decoding. By setting $r>1$, we retain more tokens than our final budget ($\tau$) allows, giving us a richer semantic pool to select from during decoding. We select the top $\lfloor r\cdot\tau\rfloor$ tokens with the highest $\alpha_{i}$ values across all sentence buckets.
@@ 5 Experiments
All the experiments are conducted using two NVIDIA H100 80GB GPUs and two NVIDIA H100 NVL 96GB GPUs. We evaluate our method on three benchmarks: PG-19 (language modeling task) [22], LongBench [2], NIAH (retrieval of a hidden “needle”) [12]. Our experiments use the models Llama-3.1-8B-Instruct [1] and Longchat-v1.5-7B [18]. We compare SentenceKV with several baselines, including H2O [29], SnapKV [14], Quest [25], as well as a Full KV cache baseline. The evaluation metrics include model accuracy, GPU memory usage (in GB), and latency (in ms/token).
@@ Results.
Table 1 reports accuracy on LongBench tasks. The full KV cache achieves the best performance but is impractical for long inputs due to high GPU memory usage. In contrast, SentenceKV achieves comparable accuracy across most tasks while operating under a strict token budget, demonstrating strong performance-memory trade-offs. It consistently matches or closely approaches full KV cache on both Llama-3.1-8B-Instruct and LongChat-v1.5-7B. Compared to other baselines, SentenceKV is more stable: H2O frequently fails with OOM errors on long inputs (marked as * in the table). SnapKV performs well but slightly lags behind SentenceKV on several tasks. Quest does not perform well on GQA-based models (including Llama-3.1). Overall, SentenceKV provides an effective solution for long-context inference with reduced memory usage.
@@
(d) SentenceKV Accuracy: 97.5%
@@
Figure 4 shows the retrieval accuracy results on the NIAH benchmark, comparing SentenceKV against H2O, SnapKV, and Quest with a token budget $\tau=128$ and context lengths up to 8000 tokens. SentenceKV achieves an outstanding accuracy of 97.5%, significantly outperforming other methods. SnapKV shows reasonable accuracy (78.2%), whereas Quest provides moderate performance (48.3%). Notably, H2O exhibits poor performance (25.2%), primarily due to its limitations in handling extensive contexts, further highlighting the advantage of SentenceKV in accurately and efficiently retrieving critical sentences within lengthy texts.

### v2

@@ Abstract
Large language models face significant computational and memory challenges when processing long contexts. During inference, efficient management of the key-value (KV) cache, which stores intermediate activations for autoregressive generation, is critical to reducing memory overhead and improving computational efficiency. Traditional token-level efficient KV caching methods overlook semantic information, treating tokens independently without considering their semantic relationships. Meanwhile, existing semantic-preserving KV cache management approaches often suffer from substantial memory usage and high time-to-first-token. To address these limitations, we propose SentenceKV, a novel sentence-level semantic KV caching approach designed to enhance inference efficiency while preserving semantic coherence. During prefilling, SentenceKV groups tokens based on sentence-level semantic similarity, compressing sentence representations into concise semantic vectors stored directly on the GPU, while individual KV pairs are offloaded to CPU. During decoding, SentenceKV generates tokens by selectively retrieving semantically relevant sentence-level KV entries, leveraging the semantic similarity between the prefilling-stage semantic vectors and decoding-stage queries. This ensures efficient and contextually accurate predictions, minimizing the loading of redundant or irrelevant data into GPU memory and significantly reducing memory overhead while maintaining stable inference latency, even for extremely long contexts. Extensive evaluations on benchmarks including PG-19, LongBench, Needle-In-A-Haystack, and RULER demonstrate that SentenceKV significantly outperforms state-of-the-art methods in both efficiency and memory usage, without compromising model accuracy.
@@ 1 Introduction
Large language models (LLMs) have achieved remarkable success across a wide range of natural language processing tasks (Min et al., 2023; Zhao et al., 2023). As LLMs are increasingly applied to complex reasoning and multi-step decision-making scenarios, the demand for longer context windows has grown significantly Chen et al. (2023). Recent advancements have extended the context window from the early GPT models to the latest iterations, such as GPT-o1 and beyond (OpenAI, 2023; Koubaa, 2023). However, this growth introduces substantial challenges in key-value (KV) cache efficiency, leading to memory bottlenecks and increased inference latency Shi et al. (2024).
@@
Due to the transformer-based architecture of LLMs Vaswani et al. (2017), the memory footprint of KV cache grows linearly with the context length, directly impacting GPU usage. For example, in the case of the LLaMA-3.1-8B model AI (2023), as shown in Appendix A.1, processing a 32k-token prompt during the decoding stage requires approximately 16 GB (using float16 precision) of GPU memory, which can be prohibitive for users without high-end hardware. Furthermore, longer prompts significantly increase computational overhead, as each additional token requires computing attention weights across the entire sequence. This quadratic attention complexity leads to substantial compute inefficiency, making efficient inference particularly challenging in resource-constrained environments.
@@
To mitigate these issues, several methods have been proposed to compress KV cache Shi et al. (2024). Among them, token eviction strategies based on fixed selection patterns or attention-weight-based pruning have shown promise (Xiao et al., 2023; Zhang et al., 2023). However, these approaches suffer from key limitations: they fail to adapt to the dynamic nature of token importance during decoding and often discard tokens permanently, neglecting their potential relevance in the future decoding steps. This static behavior can degrade the performance of the model, especially in tasks requiring complex reasoning and long-range dependencies Liu et al. (2025a).
@@
Another category of approaches, such as Quest Tang et al. (2024) and ShadowKV Sun et al. (2024), dynamically retrieve portions of KV cache from CPU-offloaded storage back to GPU memory during attention calculation. Typically, these methods segment the input into fixed-size chunks and selectively retrieve them based on query relevance at each decoding step. However, fixed-size chunking can disrupt semantic coherence by arbitrarily breaking natural semantic boundaries. Some dynamic retrieval approaches, including ClusterKV Liu et al. (2024) and SqueezeAttention Hooper et al. (2024), mitigate this issue by clustering semantically related keys into coherent chunks. Yet, these clustering-based methods introduce significant computational overhead, resulting in increased time-to-first-token (TTFT) latency.
@@
To address these limitations, we propose SentenceKV, a novel approach that dynamically manages KV cache based on sentence-level semantic information. Our key insight is that sentences inherently encapsulate richer semantic information compared to isolated tokens, making them a more effective unit for caching and retrieval. For instance, as illustrated in Figure 1, sentences from different domains tend to form distinct semantic clusters within the representation space of LLMs, a phenomenon also observed in recent literature Chen et al. (2024). We further propose a novel multi-query-aware mechanism for retrieving relevant sentence-level KV cache entries during decoding.
@@
Extensive experiments on long-context benchmarks, including LongBench Bai et al. (2023), Needle-In-A-Haystack (NIAH) Kamradt (2023), and RULER Hsieh et al. (2024) demonstrate that our approach not only improves inference efficiency but also maintains or even enhances model accuracy across diverse tasks. Quantitative analysis shows that our method offers an efficient solution for long-context processing.
@@ 2 Related Work
StreamingLLM Xiao et al. (2023) enables infinite-context handling via attention sinks and a moving window, allowing models to process extremely long sequences without growing memory footprints. SnapKV Li et al. (2025) preserves essential KV positions based on per-head attention patterns with an observation window. FastGen Ge et al. (2023) progressively drops tokens with low future impact, enhancing efficiency with minimal quality loss. PyramidInfer Yang et al. (2024) and PyramidKV Cai et al. (2024) propose a layer-wise pyramidal KV cache retention pattern, preserving more tokens in upper layers of the model to maximize memory savings without sacrificing context.
@@
Dynamic access mechanisms adaptively manage KV retention based on task-specific requirements. H2O Zhang et al. (2023) dynamically retains high-impact tokens based on accumulated attention scores. TOVA Oren et al. (2024) prunes tokens based on real-time query relevance, but risks removing context needed for future steps. The quest Tang et al. (2024) partitions KV cache into pageable memory and predicts critical pages per query, in which case fragmentation issues can arise. InfLLM Xiao et al. (2024) stores distant contexts into additional memory units and employs an efficient mechanism to lookup token-relevant units for attention computation. ShadowKV Sun et al. (2024) offloads the value cache to CPU and stores a low-rank key cache on GPU, using a KV selection strategy to reduce memory footprint and maintain high-throughput decoding. SqueezeAttention Hooper et al. (2024) optimizes cache allocation across both sequence and layer dimensions, assigning larger budgets to important layers while pruning unimportant ones. DynamicKV Zhou et al. (2024) implements task-aware adaptive retention, adjusting KV budgets per layer to maintain task-specific relevance while reducing cache size.
@@
As LLMs tackle increasingly complex tasks, optimizing KV cache at a semantic level is crucial to maintaining output coherence. ChunkKV Liu et al. (2025b) groups tokens into semantic chunks, retaining the most informative segments and discarding redundant ones to enhance long-context inference efficiency. ClusterKV Liu et al. (2024) introduces semantic clustering, which, unlike methods that permanently evict tokens or recall them based solely on textual positions, clusters tokens semantically and recalls them at the granularity of semantic clusters. Task-KV He et al. (2025) differentiates between heterogeneous and general aggregation attention heads, preserving full KV cache for critical heads while reducing it for less relevant ones. SepLLM Chen et al. (2024) introduces a data-dependent sparse attention mechanism to mitigate the excessive attention allocated to seemingly uninformative separator tokens. By compressing segment-level information into these tokens, SepLLM reduces KV cache while preserving essential content.
While chunk-based compression better preserves semantics than token-wise eviction, it may still fragment full thought units, omitting crucial context. Prior studies show that ignoring natural boundaries, such as sentence breaks, can harm language modeling performance (Dai et al., 2019; Rae et al., 2019). SentenceKV addresses this by retaining entire sentences or semantically self-contained segments, ensuring holistic context preservation. By bridging local semantic retention from ChunkKV with global coherence, SentenceKV further enhances long-context inference while maintaining minimal memory overhead.
@@ 3 Motivation
Existing KV cache compression methods typically operate at the token-level, where tokens considered less important—often determined by heuristics such as attention scores—are evicted during the prefilling stage. However, recent studies suggest that the relevance of prefilled tokens is not static throughout the decoding process Tang et al. (2024). Evicting tokens prematurely based on their initial perceived importance can degrade inference accuracy since tokens initially deemed irrelevant might later become crucial during decoding. To address this dynamic nature, recent studies propose retrieving essential prefilled tokens dynamically based on their relevance to the current decoding query. To mitigate computational overhead associated with searching across all prefilled tokens, these methods typically segment the prefilling text into fixed-size chunks, retrieving and utilizing only the most relevant chunks during decoding.
Figure 1: Similarities between sentence embeddings from different categories in the Pile datasets Gao et al. (2020).
However, these chunk-based strategies lack a principled justification for preserving semantic coherence, as fixed-size chunks often fragment semantic content arbitrarily. Intuitively, sentences encapsulate complete semantic units, making them natural candidates for maintaining context coherence. To empirically validate this intuition, we examined sentences from distinct categories in the Pile dataset Gao et al. (2020) and measured their embedding similarities in LLMs. As illustrated in Figure 1, sentences from different semantic categories exhibit clearly distinguishable embedding patterns. Moreover, during the decoding phase, we observed that the model-generated query tends to focus attention primarily on semantically related sentences rather than arbitrary token segments. This observation motivates our sentence-level caching approach, which segments the prefilling text into semantically coherent sentence units. During decoding, we leverage the sentence currently being generated to dynamically identify and recall the most semantically relevant sentence-level cache entries, thereby efficiently preserving context and enhancing inference accuracy.
@@ 4.1 Prefilling Stage
We first split the input text into sentences according to punctuation, creating multiple sentence buckets, each containing a varying number of tokens. For example, a 32K token input might be split into hundreds of sentence buckets of different lengths. While token importance changes dynamically during decoding, maintaining all tokens at full precision in GPU memory is inefficient and unnecessary. To determine which tokens to retain, we apply an observation window Li et al. (2025) for token importance calculation as follows:
@@
For an input prompt of length $L$ (e.g., $L=32K$), we designate the last $N$ tokens (typically $N=32$) as our observation window. We perform a forward pass where each token in positions (LN+1)(L-N+1) to $L$ calculates attention scores with respect to all preceding tokens in positions $1$ to (LN)(L-N). For each token $i$ in the preceding positions, we calculate its importance score $\alpha_{i}$ by summing the attention scores it receives from all tokens in the observation window, across all the attention heads.
@@
We set a token budget $\tau$ (e.g., $\tau=1024$) that represents the maximum number of tokens we can keep in GPU memory during decoding. Since tokens that appear unimportant during prefilling may become relevant at later decoding steps, we introduce the semantic keeping factor $r$ (typically $r=2$ or $3$), which controls how many tokens are initially retained before selective pruning during decoding. By setting $r>1$, we retain more tokens than our final budget ($\tau$) allows, giving us a richer semantic pool to select from during decoding. We select the top $\lfloor r\cdot\tau\rfloor$ tokens with the highest $\alpha_{i}$ values across all sentence buckets. This design ensures that the method works when sentence lengths are unbalanced; see Appendix A.7 for the analysis.
@@ 5 Experiments
All the experiments are conducted using two NVIDIA H100 80GB GPUs and two NVIDIA H100 NVL 96GB GPUs. We evaluate our method on four benchmarks: PG-19 (language modeling task) Rae et al. (2019), LongBench Bai et al. (2023), NIAH (retrieval of a hidden “needle”) Kamradt (2023), RULER Hsieh et al. (2024). Our experiments use the models Llama-3-8B AI (2024a), Llama-3.1-8B-Instruct AI (2023), Llama-3.2-3B-Instruct AI (2024b) and Longchat-v1.5-7B LMSYS (2024). We compare SentenceKV with several baselines, including H2O Zhang et al. (2023), SnapKV Li et al. (2025), Quest Tang et al. (2024), InfLLM Xiao et al. (2024), ShadowKV Sun et al. (2024), as well as a Full KV cache baseline. The evaluation metrics include model accuracy, GPU memory usage (in GB), and latency (in ms/token).
@@ 5.2 LongBench
Llama-3.2-3B
Full KV 23.53 41.69 49.82 49.96 41.8 20.96 2.0 9.17 5.78 7.5 78.5 26.31 25.2
H2O * * * * * * * * * * * * *
SnapKV 20.02 35.17 48.29 48.98 37.63 20.12 0.5 9.07 6.19 4.35 81.5 26.19 25.36
Quest 15.53 31.7 41.36 46.07 33.25 19.12 3.5 13.17 8.36 5.17 67.0 28.64 28.45
InfLLM 20.48 33.97 45.89 45.94 36.7 16.54 0.0 48.77 3.56 2.0 36.5 24.86 26.27
SentenceKV 24.12 38.38 49.38 48.98 40.01 20.95 6.0 8.82 6.06 0.0 84.0 26.64 23.46
@@ Results.
Table 1 reports accuracy on LongBench tasks. The full KV cache achieves the best performance but is impractical for long inputs due to high GPU memory usage. In contrast, SentenceKV achieves comparable accuracy across most tasks while operating under a strict token budget, demonstrating strong performance-memory trade-offs. It consistently matches or closely approaches full KV cache on both Llama models and LongChat-v1.5-7B. Compared to other baselines, SentenceKV is more stable: H2O frequently fails with OOM errors on long inputs (marked as * in the table). SnapKV and InfLLMperforms well but slightly lags behind SentenceKV on several tasks. Quest does not perform well on GQA-based models (including Llama-3.1). Overall, SentenceKV provides an effective solution for long-context inference with reduced memory usage.
@@
(d) InfLLM Accuracy: 87.8%
(e) ShadowKV Accuracy: 85.8%
(f) SentenceKV Accuracy: 97.5%
@@
Figure 4 shows the retrieval accuracy results on the NIAH benchmark, comparing SentenceKV against H2O, SnapKV, Quest, InfLLMand ShadowKVwith a token budget $\tau=128$ and context lengths up to 8000 tokens, with Llama-3-8B model. SentenceKV achieves an outstanding accuracy of 97.5%, significantly outperforming other methods. SnapKV, InfLLM and ShadowKV shows reasonable accuracy (78.2%, 87.8% and 85.8%), whereas Quest provides moderate performance (48.3%). The reason for Quest’s low accuracy is analyzed in Appendix A.4. Notably, H2O exhibits poor performance (25.2%), primarily due to its limitations in handling extensive contexts, further highlighting the advantage of SentenceKV in accurately and efficiently retrieving critical sentences within lengthy texts.
5.4 RULER
Setup.
Following the LongBench and NIAH evaluations, we additionally benchmark SentenceKV on the RULER long-context retrieval suite. Each example contains up to $64\mathrm{K}$ tokens, pushing GPU memory pressure well beyond the typical setting. We use Llama-3.1-8B-Instruct with flash–attention enabled and compare against representative KV-cache baselines.
Results.
Table 2 reports retrieval accuracy (higher is better) on eight RULER subtasks. SentenceKV matches or exceeds every offloading baseline and trails the full kv cache upper-bound on specific tasks. ShadowKV attains comparable accuracy, but incurs longer prefill time owing to its per-sentence SVD reconstruction; in contrast, SentenceKV keeps the pipeline simple, light-weight, and easier to integrate.
Method S1 S2 MK1 MQ MV FWE VT QA1
FullKV 1.00 1.00 1.00 0.98 0.96 0.85 0.94 0.55
SnapKV 1.00 0.94 0.96 0.82 0.76 0.55 0.82 0.63
H2O OOM OOM OOM OOM OOM OOM OOM OOM
Quest 1.00 0.98 1.00 0.82 0.87 0.61 0.74 0.56
InfLLM 1.00 0.20 0.14 0.17 0.22 0.81 0.43 0.12
ShadowKV 1.00 0.98 1.00 0.87 0.74 0.78 0.79 0.70
SentenceKV 1.00 0.98 0.98 0.77 0.78 0.61 0.86 0.63
Table 2: Retrieval accuracy on the RULER benchmark ($64\mathrm{K}$ tokens). “OOM” indicates out-of-memory on a single H100 80 GB GPU.
Although a fixed semantic keeping factor already performs well, we notice task-dependent variation. An adaptively chosen semantic keeping factor may increase the accuracy; we leave a full exploration of this adaptive strategy to future work. Overall, these RULER results confirm the scalability and robustness of SentenceKV in extreme long-context scenarios and motivate its deployment in real-world applications where very long contexts are becoming the norm.